OpenStudy (anonymous):

I'm having trouble solving this please help! Properties of Exponents (-2x^3)^3 multiplied by (-5x^4)^2

OpenStudy (robertok):

Here you go Simplifying (1 + x)(3 + 2x) Multiply (1 + x) * (3 + 2x) (1(3 + 2x) + x(3 + 2x)) ((3 * 1 + 2x * 1) + x(3 + 2x)) ((3 + 2x) + x(3 + 2x)) (3 + 2x + (3 * x + 2x * x)) (3 + 2x + (3x + 2x2)) Combine like terms: 2x + 3x = 5x (3 + 5x + 2x2)

jhonyy9 (jhonyy9):

-2^3 = -8 (x^3)^3 = x^9

jhonyy9 (jhonyy9):

-5^2 =25 (x^4)^2 = x^8

OpenStudy (anonymous):

Okay, thank you both! It makes more sense now.

OpenStudy (robertok):

your welcome i hope what i said helped

OpenStudy (anonymous):

It did!

OpenStudy (robertok):

lol but why didnt i get a medal for my comments lol

OpenStudy (anonymous):

Happy now?

OpenStudy (robertok):

lol yes xD i have only had 1 medal before xD and if you need help just ask im really great at math im doing pre cal

OpenStudy (anonymous):

Okay will do. Thanks again :)

OpenStudy (robertok):

welcome :)